# Controlling Continuous Relaxation for Combinatorial Optimization

### v3

@@ 5.3 Diverse bipartite matching
To evaluate the applicability of the CRA-PI-GNN solver to more practical problems not on graphs, we conducted experiments on DBM problems (Ferber et al., 2020; Mulamba et al., 2020; Mandi et al., 2022); refer to Appendix E.1 for details.
This problems consists of $27$ distinct instances with varying properties, and each instance comprises $100$ nodes representing scientific publications, divided into two groups of $50$ nodes $N_{1}$ and $N_{2}$. The optimization is formulated as follows:
@@ 6 Conclusion
This study proposes CRA strategy to address the both optimization issue and rounding issue in UL-based solvers. The CRA strategy introduces a penalty term that dynamically shifts from prioritizing continuous solutions where the non-convexity of the objective function are effectively smoothed, to enforcing discreteness, eliminating the artificial rounding. Experimental results demonstrate that CRA-PI-GNN solver significantly outperforms PI-GNN solver and greedy algorithms across various complex CO problems, including MIS, MaxCut, and DBM problems. CRA approach not only enhances quality of solutions but also accelerates the learning process.
@@ E.2 GNNs
Appendix F Additional experiments
In this sectioin, we will examine the issue (I) with continuous relaxations and the issue (II), the difficulties of optimization, as pointed out by previous studies [Wang and Li, 2023, Angelini and Ricci-Tersenghi, 2023], in the NP-hard problems of MIS and the MaxCut problem.
Therefere, we conducted numerical experiments using the PI-GNN solver for MIS and MaxCut problems on RRGs with higher degrees. To ensure the experimental impartiality, we adhered to the original settings of the PI-GNN solver [Schuetz et al., 2022b]. Refer to Section E for the detailed experimental settings. Fig. 7 (top) shows the solutions obtained by the PI-GNN solver as a function of the degree $d$ for the MIS and MaxCut problems with varying system sizes $N$. These results indicate that finding independent and cut sets becomes unfeasible as the RRG becomes denser. In addition, to clarify the reasons for these failures, we analyzed the dynamics of the cost function for MIS problems with $N=10{,}000$, with a specific focus on a graph with degrees $d=5$ and $d=20$, as depicted in Fig.  7 (bottom). For the $d=5$ case, the cost function goes over the plateau of $\hat{l}({\bm{\theta}};G,{\bm{\lambda}})=0$ with ${\bm{p}}_{\theta}(G)={\bm{0}}_{N}$, as investigated in the histogram, eventually yielding a solution comparable to those presented by Schuetz et al. [2022a]. Conversely, in the $d=20$ case, the cost function remains stagnant on the plateau of $\hat{l}({\bm{\theta}};G,{\bm{\lambda}})=0$ with ${\bm{p}}_{\theta}(G)={\bm{0}}_{N}$, thereby failing to find any independent nodes. Interpreting this phenomenon, we hypothesize that the representation capacity of the GNN is sufficiently large, leading us to consider the optimization of $\hat{L}_{\mathrm{MIS}}({\bm{\theta}};G,\lambda)$ and $\hat{L}_{\mathrm{MaxCut}}({\bm{\theta}};G)$ as a variational optimization problem relative to ${\bm{p}}_{\theta}$. In this case, ${\bm{p}}_{\theta}^{\ast}={\bm{0}}_{N}$ satisfies the first-order variational optimality conditions $\delta\hat{l}_{\mathrm{MIS}}/\delta{\bm{p}}_{\theta}|_{{\bm{p}}_{\theta}={\bm{p}}^{\ast}}=\delta\hat{l}_{\mathrm{MaxCut}}/\delta{\bm{p}}_{\theta}|_{{\bm{p}_{\theta}}={\bm{p}}^{\ast}}={\bm{0}}_{N}$, which implies a potential reason for absorption into the plateau. However, this does not reveal the conditions for the convergence to the fixed point ${\bm{p}}^{\ast}$ during the early learning stage or the condition to escape from the fixed point ${\bm{p}}^{\ast}$. Thus, an extensive theoretical evaluation through stability analysis remains an important topic for future work.
F.2 Additional results of Gset
F.3 Ablation over initial scheduling value and scheduling rate
We conducted an ablation study focusing on the initial scheduling value $\gamma(0)$ and scheduling rate $\varepsilon$. This numerical experiment was conducted under the configuration described in Section 5 and E. Fig.  8 shows the IS density of $N=10000$ MIS problems on a $100$-RRG as a function of the initial scheduling value $\gamma(0)$ and the scheduling rate $\varepsilon$ using the CRA-PI-GNN with both GraphSage and GCV. As can be seen, smaller initial scheduling $\gamma(0)$ and scheduling rate $\varepsilon$ values typically yield better solutions. However, the convergence time increases progressively as the initial scheduling $\gamma(0)$ and scheduling rate $\varepsilon$ values become smaller. In addition, GraphSage consistently produces better solutions even with relatively larger initial scheduling $\gamma(0)$ and scheduling rate $\varepsilon$ values, which implies that the GNN architecture influences both the solution quality and the effective regions of the initial scheduling $\gamma(0)$ and scheduling rate $\varepsilon$ values for the annealing process.
@@ F.3 Ablation over initial scheduling value and scheduling rate
F.4 Ablation over curve rate
@@
Figure 9: (Left) Independent set density as a function of curveture rate $\alpha$ in Eq. (5). (Right) Cut ratio as a function of curveture rate $\alpha$ in Eq. (5). Error bars represent the standard deviations of the results

### v4

@@ 5.4 Diverse bipartite matching
To evaluate the applicability of the CRA-PI-GNN solver to more practical problems not on graphs, we conducted experiments on DBM problems (Ferber et al., 2020; Mulamba et al., 2020; Mandi et al., 2022); refer to Appendix E.1 for details. This problems consists of $27$ distinct instances with varying properties, and each instance comprises $100$ nodes representing scientific publications, divided into two groups of $50$ nodes $N_{1}$ and $N_{2}$. The optimization is formulated as follows:
@@ 6 Conclusion
This study proposes CRA strategy to address the both optimization and rounding issue in UL-based solvers. CRA strategy introduces a penalty term that dynamically shifts from prioritizing continuous solutions, where the non-convexity of the objective function is effectively smoothed, to enforcing discreteness, thereby eliminating artificial rounding. Experimental results demonstrate that CRA-PI-GNN solver significantly outperforms both PI-GNN solver and greedy algorithms across various complex CO problems, including MIS, MaxCut, and DBM problems. CRA approach not only enhances solution quality but also accelerates the learning process.
Limitation
In this numerical experiments, most hyperparameters were fixed to their default values, as outlined in Section 5.1, with minimal tuning. However, tuning may be necessary for certain problems or to further enhance performance.
@@ E.2 GNNs
Appendix F Additional experiments
In this sectioin, we will examine the issue (I) with continuous relaxations and the issue (II), the difficulties of optimization, as pointed out by previous studies [Wang and Li, 2023, Angelini and Ricci-Tersenghi, 2023], in the NP-hard problems of MIS and the MaxCut problem. Therefere, we conducted numerical experiments using the PI-GNN solver for MIS and MaxCut problems on RRGs with higher degrees. To ensure the experimental impartiality, we adhered to the original settings of the PI-GNN solver [Schuetz et al., 2022b]. Refer to Section E for the detailed experimental settings. Fig. 7 (top) shows the solutions obtained by the PI-GNN solver as a function of the degree $d$ for the MIS and MaxCut problems with varying system sizes $N$. These results indicate that finding independent and cut sets becomes unfeasible as the RRG becomes denser. In addition, to clarify the reasons for these failures, we analyzed the dynamics of the cost function for MIS problems with $N=10{,}000$, with a specific focus on a graph with degrees $d=5$ and $d=20$, as depicted in Fig.  7 (bottom). For the $d=5$ case, the cost function goes over the plateau of $\hat{l}({\bm{\theta}};G,{\bm{\lambda}})=0$ with ${\bm{p}}_{\theta}(G)={\bm{0}}_{N}$, as investigated in the histogram, eventually yielding a solution comparable to those presented by Schuetz et al. [2022a]. Conversely, in the $d=20$ case, the cost function remains stagnant on the plateau of $\hat{l}({\bm{\theta}};G,{\bm{\lambda}})=0$ with ${\bm{p}}_{\theta}(G)={\bm{0}}_{N}$, thereby failing to find any independent nodes. Interpreting this phenomenon, we hypothesize that the representation capacity of the GNN is sufficiently large, leading us to consider the optimization of $\hat{L}_{\mathrm{MIS}}({\bm{\theta}};G,\lambda)$ and $\hat{L}_{\mathrm{MaxCut}}({\bm{\theta}};G)$ as a variational optimization problem relative to ${\bm{p}}_{\theta}$. In this case, ${\bm{p}}_{\theta}^{\ast}={\bm{0}}_{N}$ satisfies the first-order variational optimality conditions $\delta\hat{l}_{\mathrm{MIS}}/\delta{\bm{p}}_{\theta}|_{{\bm{p}}_{\theta}={\bm{p}}^{\ast}}=\delta\hat{l}_{\mathrm{MaxCut}}/\delta{\bm{p}}_{\theta}|_{{\bm{p}_{\theta}}={\bm{p}}^{\ast}}={\bm{0}}_{N}$, which implies a potential reason for absorption into the plateau. However, this does not reveal the conditions for the convergence to the fixed point ${\bm{p}}^{\ast}$ during the early learning stage or the condition to escape from the fixed point ${\bm{p}}^{\ast}$. Thus, an extensive theoretical evaluation through stability analysis remains an important topic for future work.
F.2 Additional results of MIS
Table 3: ApR and runtime are evaluated on three benchmarks provided by DIMES [Qiu et al., 2022]. The ApR is assessed relative to the results obtained by KaMIS. Runtime is reported as the total clock time, denoted in seconds (s), minutes (m), or hours (h). The runtime and solution quality are sourced from iSCO [Sun et al., 2023]. The baselines include solvers from the Operations Research (OR) community, as well as data-driven approaches utilizing Reinforcement Learning (RL), Supervised Learning (SL) combined with Tree Search (TS), Greedy decoding (G), or sampling (S). Methods that fail to produce results within 10 times the time limit of DIMES are marked as N/A.
Method Type ER-[700-800] ER-[9000-11000]
ApR Time ApR Time
KaMIS OR 1.000 52.13m 1.000 7.6h
Gurobi OR 0.922 50.00m N/A N/A
Intel SL+TS 0.865 20.00m N/A N/A
SL+G 0.777 6.06m 0.746 5.02m
DGL SL+TS 0.830 22.71m N/A N/A
LwD RL+S 0.918 6.33m 0.907 7.56m
DIMES RL+G 0.852 6.12m 0.841 5.21m
RL+S 0.937 12.01m 0.873 12.51m
iSCO fewer steps 0.998 1.38m 0.990 9.38m
more steps 1.006 5.56m 1.008 1.25h
CRA UL-based 0.928 47.30m 0.963 1.03h
We evaluate our method using the MIS benchmark dataset from recent studies [Goshvadi et al., 2023, Qiu et al., 2022], which includes graphs from SATLIB [Hoos and Stützle, 2000] and Erdős–Rényi graphs (ERGs) of varying sizes. Following Sun et al. [2023], our test set consists of $500$ SATLIB graphs, each containing between $403$ and $449$ clauses with up to $1{,}347$ nodes and $5{,}978$ edges, $128$ ERGs with $700$ to $800$ nodes each, and $16$ ERGs with $9{,}000$ to $11{,}000$ nodes each. We conducted numerical experiments on PQQA using four different configurations: parallel runs with $S=100$ or $S=1000$ and shorter steps (3000 steps) or longer steps (30000 steps), similar to the approach in iSCO [Sun et al., 2023]. Table 3 presents the solution quality and runtime results. The results show that CRA, which optimizes the relaxed variables as an optimization of GNN parameters, takes extra time for smaller ER-[700-800] instances due to the smaller number of decision variables. However, for larger instances, CRA achieves results comparable to iSCO. Although limited space makes it difficult to present other benchmark results employed by iSCO, such as MaxCut and MaxClique, numerical experiments on these benchmarks also show that CRA is less effective for small problems. However, for larger problems, the results are comparable to or slightly inferior to those of iSCO.
We also investigated the relationship between the order of the graph and the solving time of the solver, and the results are shown in Table 4 and 5. The results demonstrate that the runtime remains nearly constant as graph order and density increase, indicating effective scalability with denser graphs.
Table 4: ApR of the MIS problem on $\mathrm{RRG}(N,d)$. All the results are averaged based on 5 $\mathrm{RRG}$s with different random seeds.
Problem ApR (CRA) ApR (PI) Time (CRA) Time (PI)
$\mathrm{RRG}(1{,}000,10)$ 0.95 0.78 108 (s) 98 (s)
$\mathrm{RRG}(1{,}000,20)$ 0.95 0.56 103 (s) 92 (s)
$\mathrm{RRG}(1{,}000,30)$ 0.94 0.00 102 (s) 88 (s)
$\mathrm{RRG}(1{,}000,40)$ 0.93 0.00 101 (s) 82 (s)
$\mathrm{RRG}(1{,}000,50)$ 0.92 0.00 102 (s) 82 (s)
$\mathrm{RRG}(1{,}000,60)$ 0.91 0.00 101 (s) 91 (s)
$\mathrm{RRG}(1{,}000,70)$ 0.91 0.00 101 (s) 86 (s)
$\mathrm{RRG}(1{,}000,80)$ 0.91 0.00 102 (s) 93 (s)
$\mathrm{RRG}(5{,}000,10)$ 0.93 0.77 436 (s) 287 (s)
$\mathrm{RRG}(5{,}000,20)$ 0.95 0.74 413 (s) 280 (s)
$\mathrm{RRG}(5{,}000,30)$ 0.95 0.00 419 (s) 283 (s)
$\mathrm{RRG}(5{,}000,40)$ 0.94 0.00 429 (s) 293 (s)
$\mathrm{RRG}(5{,}000,50)$ 0.94 0.00 418 (s) 324 (s)
$\mathrm{RRG}(5{,}000,60)$ 0.93 0.00 321 (s) 302 (s)
$\mathrm{RRG}(5{,}000,70)$ 0.92 0.00 321 (s) 325 (s)
$\mathrm{RRG}(5{,}000,80)$ 0.92 0.00 330 (s) 305 (s)
Table 5: The ApR of the MIS problem on $\mathrm{ERG}(N,d)$ is evaluated against the results of KaMIS. Due to time limitations, the maximum running time for KaMIS was constrained. The results below present the average ApRs and runtimes across five different random seeds.
Problem CRA (ApR) PI (ApR) Time (CRA) Time (PI) Time (KaMIS)
$\mathrm{ERG}(1{,}000,0.05)$ 0.97 0.01 103 (s) 98 (s) 100 (s)
$\mathrm{ERG}(1{,}000,0.10)$ 0.95 0.00 100 (s) 98 (s) 210 (s)
$\mathrm{ERG}(1{,}000,0.15)$ 0.94 0.00 100 (s) 92 (s) 315 (s)
$\mathrm{ERG}(1{,}000,0.20)$ 0.91 0.00 99 (s) 88 (s) 557 (s)
$\mathrm{ERG}(1{,}000,0.25)$ 0.93 0.00 98 (s) 82 (s) 733 (s)
$\mathrm{ERG}(1{,}000,0.30)$ 0.90 0.00 98 (s) 82 (s) 1000 (s)
$\mathrm{ERG}(1{,}000,0.35)$ 0.92 0.00 99 (s) 91 (s) 1000 (s)
$\mathrm{ERG}(1{,}000,0.40)$ 0.91 0.00 97 (s) 86 (s) 1000 (s)
F.3 Additional results of Gset
@@
F.4 Additional results of TSP
We conducted additional experiments on several TSP problems from the TSPLIB dataset [^1] http://comopt.ifi.uni-heidelberg.de/software/TSPLIB95/, presenting results that illustrate the $\alpha$-dependency.
Experiments calculated the ApR as the ratio of the optimal value to the CRA result, with the ApR representing the average and standard deviation over 3 seeds. The “–” symbol in PI-GNN denotes cases where most variables are continuous values and where no solution satisfying the constraint was found within the maximum number of epochs. The same GNN and optimizer settings were used as in the main text experiments in Section 5.1.
Table 7 shows that the CRA approach yielded solutions with an ApR exceeding 0.9 across various instances. Notably, for the burma14 problem, our method identified the global optimal solution (3,323) multiple times. However, the optimal value can vary based on the specific GNN architecture and problem structure, indicating that a more comprehensive ablation study could provide valuable insights in future work.
@@ F.4 Additional results of TSP
Figure 9: (Left) Independent set density as a function of curveture rate $\alpha$ in Eq. (5). (Right) Cut ratio as a function of curveture rate $\alpha$ in Eq. (5). Error bars represent the standard deviations of the results.
Table 7: Comparison of ApR performance across different values of $p$ for various TSPLIB instances.
burma14 ulysses22 st70 gr96
ApR ($p=2$) $0.91\pm 0.08$ $0.89\pm 0.03$ $0.96\pm 0.01$ $0.81\pm 0.05$
ApR ($p=4$) $0.98\pm 0.10$ $0.92\pm 0.02$ $0.85\pm 0.03$ $0.82\pm 0.03$
ApR ($p=6$) $0.97\pm 0.14$ $0.88\pm 0.07$ $0.88\pm 0.04$ $0.90\pm 0.05$
ApR ($p=8$) $0.99\pm 0.06$ $0.89\pm 0.05$ $0.80\pm 0.02$ $0.86\pm 0.05$
ApR (PI) $0.736\pm 1.21$
Optimum 3,323 7,013 675 5,5209
F.5 Ablation over initial scheduling value and scheduling rate
We conducted an ablation study focusing on the initial scheduling value $\gamma(0)$ and scheduling rate $\varepsilon$. This numerical experiment was conducted under the configuration described in Section 5 and E. Fig.  8 shows the IS density of $N=10{,}000$ for MIS problems on a $100$-RRG as a function of the initial scheduling value $\gamma(0)$ and the scheduling rate $\varepsilon$ using the CRA-PI-GNN with both GraphSage and GCV. As can be seen, smaller initial scheduling $\gamma(0)$ and scheduling rate $\varepsilon$ values typically yield better solutions. However, the convergence time increases progressively as the initial scheduling $\gamma(0)$ and scheduling rate $\varepsilon$ values become smaller. In addition, GraphSage consistently produces better solutions even with relatively larger initial scheduling $\gamma(0)$ and scheduling rate $\varepsilon$ values, which implies that the GNN architecture influences both the solution quality and the effective regions of the initial scheduling $\gamma(0)$ and scheduling rate $\varepsilon$ values for the annealing process.
F.6 Ablation over curve rate